OpenStudy (anonymous):

The value of y varies directly with x, and y=-18 when x=6. Find y when x=-8 Fink k: Use k to find y: y=kx y=(____) (____) ____=k y=____________

OpenStudy (anonymous):

The value of y varies directly with x, and y= 1/2 when x=5. Find y when x=30. Find k: Use k to find y: y=kx y= (_____) (_____) ______=k y=_______

jimthompson5910 (jim_thompson5910):

Are these 2 different problems?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

Please post one problem at a time

jimthompson5910 (jim_thompson5910):

Which one did you need help with?

OpenStudy (anonymous):

The 1st one

jimthompson5910 (jim_thompson5910):

The value of y varies directly with x, and y=-18 when x=6. Find y when x=-8 Fink k: Use k to find y: y=kx y=(____) (____) ____=k y=____________

OpenStudy (anonymous):

What does k stand for?

jimthompson5910 (jim_thompson5910):

y=kx -18=k*(6) ... plug in y=-18 and x=6 -18/6 = k -3 = k k = -3 and k is just an unknown constant

jimthompson5910 (jim_thompson5910):

So the equation is y = -3x

OpenStudy (anonymous):

Ohhhhhh ok thanks!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (mertsj):

What other 2 are you talking about?